# Pure Vision Language Action (VLA) Models: A Comprehensive Survey

### v2

@@ 7.1.5 Evaluation and Benchmarking Limitations
7.2 Opportunities of Vision Language Action Models
@@ 8 Conclusion
Recent advances in Vision Language Action (VLA) models have extended the generalizable capabilities of vision language models to robotic applications, including embodied AI, autonomous driving, and diverse manipulation tasks. This survey systematically charts the emergence of VLA approaches by examining their motivations, methodologies, and applications. It provides a unified taxonomy of architectural structures and analyzes over 300 articles, along with supporting materials. We first categorize VLA architectural innovations based on autoregression-based models, diffusion-based models, reinforcement-based learning, hybrid structures, and efficiency optimization techniques. Subsequently, we explore the datasets, benchmarks, and simulation platforms that facilitate VLA training and evaluation. Building on this comprehensive review, we analyze the strengths and limitations of current approaches and highlight potential directions for future research. Collectively, these insights offer a consolidated reference and a forward-looking roadmap for developing trustworthy, continually evolving VLAs capable of advancing general artificial intelligence in robotic systems.

### v3

@@ 7.1.5 Evaluation and Benchmarking Limitations
7.2 Opportunities of Vision-Language-Action Models
@@ 8 Conclusion
Recent advances in Vision-Language-Action (VLA) models have extended the generalizable capabilities of vision-language models to robotic applications, including embodied AI, autonomous driving, and diverse manipulation tasks. This survey systematically charts the emergence of VLA approaches by examining their motivations, methodologies, and applications. It provides a unified taxonomy of architectural structures and analyzes over 300 articles, along with supporting materials. We first categorize VLA architectural innovations based on autoregression-based models, diffusion-based models, reinforcement-based learning, hybrid structures, and efficiency optimization techniques. Subsequently, we explore the datasets, benchmarks, and simulation platforms that facilitate VLA training and evaluation. Building on this comprehensive review, we analyze the strengths and limitations of current approaches and highlight potential directions for future research. Collectively, these insights offer a consolidated reference and a forward-looking roadmap for developing trustworthy, continually evolving VLAs capable of advancing general artificial intelligence in robotic systems.